# Peculiar dark matter halos inferred from gravitational lensing as a manifestation of modified gravity

### v2

@@ 6.3 Pair of point masses observed aligned
Surface density. The results for surface density are presented in Figs. 25 and 26. The profile shows, similarly to the case of a single point mass, a peak near the MOND transitional radius, forming a ring when observed on the sky (Milgrom & Sanders, 2008). The surface density of the PDM near the ring is about $10^{15}\,M_{\sun}\,$pc-2. The height of the peak grows with the $\eta$ parameter, since the total mass of the system increases. Increasing the $\theta$ parameter (i.e. decreasing the separation between the point masses while keeping their masses fixed), leads to a decrease of the surface density at a given distance from the origin [^2] It might be confusing that $\Sigma_{\mathrm{ph}}$ here decreases with increasing $\theta$, while it was increasing for the perpendicular view in the Fig. 14, where we showed that this is a universal consequence of Eq. 4. The contradiction is just apparent: in Fig. 14, we plotted a spatial range that is fixed in the units of the distance between the points, while the range plotted in Fig. 24 is fixed in the units of kpc or arcsec. In other words, in Fig. 14, with increasing $\theta$, we were plotting smaller and smaller region around the system center, where the density is the highest. If the profiles of $\Sigma_{\mathrm{ph}}$ in Fig. 24 are plotted with the radial distance expressed in the units of $d$, then indeed $\Sigma_{\mathrm{ph}}$ increases with $\theta$ at fixed $r/d$. . This is because the gravitational field of one point mass is progressively more reduced by the EFE coming from the other point mass, as $\theta$ increases, such that the density of the PDM decreases. If the point masses are far from each other, such that they reduce the gravity of each other the least, then Eq. 22 tells us that the surface density is proportional to $\sqrt{M_{1}}+\sqrt{M_{2}}$. In the opposite limit, when the two point masses coincide, the same equation indicates that the surface density is proportional to $\sqrt{M_{1}+M_{2}}$, that is lower. The strength of the external field again only has a negligible impact on the appearance of the inner profile of the PDM halo.
@@
Surface density difference. The difference of the surface densities of PDM and the superposition density is generally on the order $10^{-13}\,M_{\sun}\,$pc-2. Surprisingly, this is rather universal value valid for all radial distances from the symmetry center and for most of the explored models. The shapes of the profiles beyond about $10^{0}-10^{1}$ kpc are caused by the same effects as discussed for the perpendicular configuration. Near the center of symmetry, the superposition surface density can be both higher or lower than the superposition surface density. To see why, we note in Fig. 2 for a single point mass, that the density of the PDM around the point mass is falls steeply toward zero. The surface density near the point mass is therefore determined by the density further away from the point mass. If two point masses seen along the line-of-sight are close to each other (i.e., low $\theta$), the external field of one of them reduces the volume density of the PDM of the other point mass. This phenomenon is absent for the PDM density and therefore its surface density can be higher than that of the actual PDM if the point masses are close to each other. On the other hand, if the point masses are far apart, the regions close to each of the point masses are almost equal to the superposition densities. Nevertheless, the combined gravitational fields of the two point masses resist easier to the global external field than if the point masses were isolated. Thus, in this case, the superposition PDM surface density is lower than the actual PDM surface density. The enhanced resistance of the actual PDM density to the global external field, when compared to the superposition PDM density, explains the behavior for the varying intensity of the global external field.
@@ 7 Observational aspects
In Fig. 4, we inspected the deviation of the QUMOND-predicted profile of the PDM from the profile expected in the $\Lambda$CDM cosmology for NFW halos. Actually, it has already been shown recently that the halos of isolated galaxies prefer MOND over NFW halos (Mistele et al., 2023b). It would be interesting to inspect whether the ratio of the observed surface density to the surface density of the MOND isothermal PDM halo (Eq. 22) shows the curvature because of the EFE, as seen in the fourth panel of Fig. 2.
@@
Bartelmann & Maturi (2017) state that $n=30-50$ galaxies per square arcminute for images by space-based telescopes. In the examples below, we will assume the optimist value $n=50\,$arcmin-2. The recently launched Euclid satellite is expected to have similar capacities and will cover one third of the extragalactic sky.
In Fig. 4, we see that the QUMOND tangential shear for the considered model differs from the tangential shear in $\Lambda$CDM in the radial range of ca. 4-20″. The difference is $10^{-3}$. If we use all background galaxies in such an annular aperture, we need to stack 2 400 lenses for a one sigma detection of this feature, or nine times as much for a three sigma detection.
@@ 8.1 Results for a point mass
  • We explored how the predictions of QUMOND and the $\Lambda$CDM cosmology differ on one example of a galaxy with the baryonic mass of $4\times 10^{10}\,M_{\sun}$. In $\Lambda$CDM, the halo has a NFW profile with parameters determined by the scaling relations with the stellar mass. The difference from the logarithmic QUMOND halo (Fig. 4) should be easily distinguishable by observations (Sect. 7).
@@ Appendix A General form of density and surface density of the PDM in the deep-MOND regime
where ${\alpha}$ is a vector of dimensionless parameters that describe the distribution of baryonic matter in the system and the distribution and orientation of the external field in the absence of the system in question. Let us consider a test particle at the position ${x}$ that moves only under the influence of gravity. It is subject of acceleration
${a}=\int\frac{M}{d^{3}}f\left(\frac{MG}{d^{2}a_{0}},\@vec{\alpha},\frac{\@vec{x^{\prime}}}{d}\frac{g_{e}}{a_{0}}\right)K(\@vec{x^{\prime}},\@vec{x})\,\mathrm{d}^{3}\!\@vec{x^{\prime}},\\ K(\@vec{x^{\prime}},\@vec{x})=\frac{G(\@vec{x}-\@vec{x^{\prime}})}{\left|\@vec{x}-\@vec{x^{\prime}}\right|^{3}}.$ (39)
@@
[…]
where we used the fact that the transformation changes the acceleration of the test particle to ${a}/\lambda$, just as the external field to $g_{e}/\lambda$ (because the bodies which generate the field are now further away). After substituting ${x^{\prime}}/\lambda=\@vec{y^{\prime}}$, one gets
${a}=\int\frac{M}{d^{3}}\frac{1}{\lambda}f\left(\frac{MG}{\lambda^{2}d^{2}a_{0}},\@vec{\alpha},\frac{\@vec{y^{\prime}}}{d},\frac{g_{e}}{\lambda a_{0}}\right)K(\@vec{y^{\prime}},\@vec{x})\,\mathrm{d}^{3}\!\@vec{y^{\prime}},$ (41)
@@
${a}=\int\frac{M}{d^{3}}\frac{1}{\lambda}f\left(\frac{MG}{\lambda^{2}d^{2}a_{0}},\@vec{\alpha},\frac{\@vec{y^{\prime}}}{d},\frac{g_{e}d}{\sqrt{GMa_{0}}}\right)K(\@vec{y^{\prime}},\@vec{x})\,\mathrm{d}^{3}\!\@vec{y^{\prime}}.$ (42)
@@
where ${\xi}$ denotes the position in the coordinates in the projection plane.
@@ Appendix B Surface density of the PDM near the zero-acceleration point is finite
Here we prove that the surface density of the PDM near the zero-acceleration point is finite, that is that in its vicinity, $|\Sigma_{\mathrm{ph}}(\@vec{\xi})|$ is smaller than a certain finite constant for any point in the lens plane ${\xi}$. If ${e}$ denotes a unit vector toward the observer, we have:
@@ Appendix C Variations of the fiducial model
Figure 25: Surface density of the PDM for a pair of point masses observed along their connecting line. In each column, one of the parameters of the fiducial model is varied, as indicated in the legends of the columns. The black curves always correspond to the fiducial model. Top row: Surface density in the logarithmic scale. Middle row: The same in the linear scale. Bottom row: Difference between the true surface density of the PDM and the superposition surface density. The symbol “slog10” stands for the symmetrized decadic logarithm with the linear part between $-10^{11}$ and $10^{11}$.

### v3

@@ 6.3 Pair of point masses observed aligned
Surface density. The results for surface density are presented in Figs. 25 and 26. The profile shows, similarly to the case of a single point mass, a peak near the MOND transitional radius, forming a ring when observed on the sky (Milgrom & Sanders, 2008). The surface density of the PDM near the ring is about $10^{15}\,M_{\sun}\,$Mpc-2. The height of the peak grows with the $\eta$ parameter, since the total mass of the system increases. Increasing the $\theta$ parameter (i.e. decreasing the separation between the point masses while keeping their masses fixed), leads to a decrease of the surface density at a given distance from the origin [^2] It might be confusing that $\Sigma_{\mathrm{ph}}$ here decreases with increasing $\theta$, while it was increasing for the perpendicular view in the Fig. 14, where we showed that this is a universal consequence of Eq. 4. The contradiction is just apparent: in Fig. 14, we plotted a spatial range that is fixed in the units of the distance between the points, while the range plotted in Fig. 24 is fixed in the units of kpc or arcsec. In other words, in Fig. 14, with increasing $\theta$, we were plotting smaller and smaller region around the system center, where the density is the highest. If the profiles of $\Sigma_{\mathrm{ph}}$ in Fig. 24 are plotted with the radial distance expressed in the units of $d$, then indeed $\Sigma_{\mathrm{ph}}$ increases with $\theta$ at fixed $r/d$. . This is because the gravitational field of one point mass is progressively more reduced by the EFE coming from the other point mass, as $\theta$ increases, such that the density of the PDM decreases. If the point masses are far from each other, such that they reduce the gravity of each other the least, then Eq. 22 tells us that the surface density is proportional to $\sqrt{M_{1}}+\sqrt{M_{2}}$. In the opposite limit, when the two point masses coincide, the same equation indicates that the surface density is proportional to $\sqrt{M_{1}+M_{2}}$, that is lower. The strength of the external field again only has a negligible impact on the appearance of the inner profile of the PDM halo.
@@
Surface density difference. The difference of the surface densities of PDM and the superposition density is generally on the order $10^{-13}\,M_{\sun}\,$Mpc-2. Surprisingly, this is rather universal value valid for all radial distances from the symmetry center and for most of the explored models. The shapes of the profiles beyond about $10^{0}-10^{1}$ kpc are caused by the same effects as discussed for the perpendicular configuration. Near the center of symmetry, the superposition surface density can be both higher or lower than the superposition surface density. To see why, we note in Fig. 2 for a single point mass, that the density of the PDM around the point mass is falls steeply toward zero. The surface density near the point mass is therefore determined by the density further away from the point mass. If two point masses seen along the line-of-sight are close to each other (i.e., low $\theta$), the external field of one of them reduces the volume density of the PDM of the other point mass. This phenomenon is absent for the PDM density and therefore its surface density can be higher than that of the actual PDM if the point masses are close to each other. On the other hand, if the point masses are far apart, the regions close to each of the point masses are almost equal to the superposition densities. Nevertheless, the combined gravitational fields of the two point masses resist easier to the global external field than if the point masses were isolated. Thus, in this case, the superposition PDM surface density is lower than the actual PDM surface density. The enhanced resistance of the actual PDM density to the global external field, when compared to the superposition PDM density, explains the behavior for the varying intensity of the global external field.
@@ 7 Observational aspects
[…]
@@
Bartelmann & Maturi (2017) state that n=3050n=30-50 galaxies per square arcminute for images by space-based telescopes. In the examples below, we will assume the optimist value n=50n=50\,arcmin-2. The recently launched Euclid satellite is expected to have similar capacities and will cover one third of the extragalactic sky.
[…]
@@ 8.1 Results for a point mass
[…]
@@ Appendix A General form of density and surface density of the PDM in the deep-MOND regime
where $\@vec{\alpha}$ is a vector of dimensionless parameters that describe the distribution of baryonic matter in the system and the distribution and orientation of the external field in the absence of the system in question. Let us consider a test particle at the position x\@vec{x} that moves only under the influence of gravity. It is subject of acceleration
$\@vec{a}=\int\frac{M}{d^{3}}f\left(\frac{MG}{d^{2}a_{0}},\@vec{\alpha},\frac{\@vec{x^{\prime}}}{d}\frac{g_{e}}{a_{0}}\right)K(\@vec{x^{\prime}},\@vec{x})\,\mathrm{d}^{3}\!\@vec{x^{\prime}},\\ K(\@vec{x^{\prime}},\@vec{x})=\frac{G(\@vec{x}-\@vec{x^{\prime}})}{\left|\@vec{x}-\@vec{x^{\prime}}\right|^{3}}.$ (39)
@@
$\frac{\@vec{a}(\lambda\@vec{x})}{\lambda}=\int\frac{M}{\lambda^{3}d^{3}}f\left(\frac{MG}{\lambda^{2}d^{2}a_{0}},\@vec{\alpha},\frac{\@vec{x^{\prime}}}{\lambda d},\frac{g_{e}}{\lambda a_{0}}\right)K(\@vec{x^{\prime}},\lambda\@vec{x})\,\mathrm{d}^{3}\!\@vec{x^{\prime}},$ (40)
where we used the fact that the transformation changes the acceleration of the test particle to $\@vec{a}/\lambda$, just as the external field to $g_{e}/\lambda$ (because the bodies which generate the field are now further away). After substituting $\@vec{x^{\prime}}/\lambda=\@vec{y^{\prime}}$, one gets
$\@vec{a}=\int\frac{M}{d^{3}}\frac{1}{\lambda}f\left(\frac{MG}{\lambda^{2}d^{2}a_{0}},\@vec{\alpha},\frac{\@vec{y^{\prime}}}{d},\frac{g_{e}}{\lambda a_{0}}\right)K(\@vec{y^{\prime}},\@vec{x})\,\mathrm{d}^{3}\!\@vec{y^{\prime}},$ (41)
@@
$\@vec{a}=\int\frac{M}{d^{3}}\frac{1}{\lambda}f\left(\frac{MG}{\lambda^{2}d^{2}a_{0}},\@vec{\alpha},\frac{\@vec{y^{\prime}}}{d},\frac{g_{e}d}{\sqrt{GMa_{0}}}\right)K(\@vec{y^{\prime}},\@vec{x})\,\mathrm{d}^{3}\!\@vec{y^{\prime}}.$ (42)
@@
where $\@vec{\xi}$ denotes the position in the coordinates in the projection plane.
@@ Appendix B Surface density of the PDM near the zero-acceleration point is finite
Here we prove that the surface density of the PDM near the zero-acceleration point is finite, that is that in its vicinity, $|\Sigma_{\mathrm{ph}}(\@vec{\xi})|$ is smaller than a certain finite constant for any point in the lens plane $\@vec{\xi}$. If e\@vec{e} denotes a unit vector toward the observer, we have:
@@ Appendix C Variations of the fiducial model
Figure 25: Surface density of the PDM for a pair of point masses observed along their connecting line. In each column, one of the parameters of the fiducial model is varied, as indicated in the legends of the columns. The black curves always correspond to the fiducial model. Top row: Surface density in the logarithmic scale. Middle row: The same in the linear scale. Bottom row: Difference between the true surface density of the PDM and the superposition surface density. The symbol “slog10” stands for the symmetrized decadic logarithm with the linear part between $10^{11}$ and $10^{11}$.